# KAN-powered large-target detection for automotive radar

### v2

@@ II Problem Statement
Consider a scenario with $K$ large targets with dimensions significantly larger than the signal wavelength. If the $k$th target is modeled with $I_{k}$ scatterers, the received signal is given by
@@
$\tau_{k,l}^{(i)}=\frac{2}{c}\big{(}R_{k}\pm\Delta{R_{k_{i}}}+(v_{k}+\delta v_{k_{i}})(l-1)T_{\mathrm{cri}}\big{)}$ (2)
@@ III-A Construction of the Binary hypothesis
It is well known that the down-converted receive signal $z(t,l)$ follows distinct distributions under each hypothesis: under $H_{0}$ (noise-only), it is modeled as $z(t,l)\sim\mathcal{CN}(0,\sigma^{2})$ ($\sigma$ is the standard deviation), while under $H_{1}$ (target present, modeled as Swerling-3), it follows $z(t,l)\sim\chi_{4}^{2}+\mathcal{CN}(0,\sigma^{2})$, $\chi_{4}^{2}$ denotes Chi-squared distribution of degree $4$. Post square-law detection, the RD cells $Z(f_{R},f_{D})$ under $H_{0}$ follow exponential distribution [24, 4]. However, under $H_{1}$, the distribution does not converge to a closed-form expression, necessitating a study of its empirical probability density function (pdf). Figure 2 presents histograms obtained over several min-max normalized RD segments. In Fig. 2(a), target presence causes most values to concentrate in the first bin, with minimal presence in higher bins. Fig. 2(b) shows a slower decay under the null hypothesis. While both histograms are observed to follow exponential distributions, Fig. 2(c) highlights their differing decay rates $\lambda_{0}\neq\lambda_{1}$. Thus, the binary hypothesis problem is formulated as:
@@ IV-A Field data and Simulated data
The Medium Range Radar (MRR) specifications in Table I apply to both simulated and field data. Field data was collected using TI AWR2243 radar with a DCA1000 data capture card. A $4$m-long hatchback car is used for field acquisitions in scenarios that include the vehicle approaching and moving away with and without acceleration. For the specifications in Table I, the MRR provides a resolution of 0.35160.35160.3516~{}0.3516m and 0.30440.30440.3044~{}0.3044m///s in range and velocity, respectively. We construct RD segment with $17\times 7$ RD cells that correspond to (5.985.985.98~{}5.98m×$\times 2.13~{}$m///s). For a fair comparison with OS-CFAR, the reference cells used for threshold computation will be from the same RD segment excluding the cell-under-test.
For the generation of simulated data at signal-to-noise ratios (SNR) from $-25$dB to $25$dB, the Swerling-3 target model ($\chi_{4}^{2}$ distribution) is employed with additive white gaussian noise. RCS values range from $19$ to $22~{}\mathrm{dBm}^{2}$ (side views), $8.7$ to $20.5~{}\mathrm{dBm}^{2}$(front views), and $14.4$ to $24.6~{}\mathrm{dBm}^{2}$ (rear views) [27], covering cars, SUVs, and trucks at random distances within MRR.
TABLE I: Specification of FMCW radar
$f_{0}$ FMCW slope $T_{cri}$ $f_{s}$ Samples Nchirps
@@ IV-B1 KAN for hypothesis classification
Unlike conventional evaluation of detection techniques, large-target detection has to be evaluated within the entire detection pipeline. Detection accuracy is defined as the percentage of detections that lie within the target RD spread. Before comparing with the existing CFAR techniques adopted in the automotive context, the proposed KAN-based detection technique is evaluated against the number of histogram bins and the testing dataset. KAN developed for $M=5$ and $M=10$ histogram bins are studied for their accuracies in Table II for (A) trained and tested on simulated data, (B) trained on simulated data, tested directly on field data, (C) trained on simulated data, fine-tuned with $14$ field data samples, and tested on field data. The KAN with MultKAN feature trained, has the specifications and parameters listed in Table. III.
TABLE II: KAN symbolic expression performance
@@
TABLE III: KAN Model specifications and parameters
Width Grids spline Optimization,Loss Prune(Node:Edge)
[$M$,2] 3 3 LBFGS,Crossentropy 0.01:0.03
From Table II, it is evident that the expressions derived from simulated data show moderate detection performance on field data (Scenario (B)). However, learning the KAN expression with as few as 14 samples in Scenario (C) significantly enhances detection performance, emphasizing the importance of transfer learning. The symbolic expressions thus yielded by KAN for $M=10$ and $M=5$ are
@@
One of these expressions is integrated with the RSP pipeline shown in Fig. 4. Notably, the KAN expression relies on only the first histogram bin $x_{0}$ for its decision in both expressions.
@@ IV-B2 Performance of detection pipeline
The proposed KAN-powered technique is compared with OS-CFAR with reference window matching the RD-segment and threshold design based on $P_{\mathrm{FA}}\in[10^{-3},10^{-6}]$ using both simulated and field data. The $P_{\mathrm{D}}$ and $P_{\mathrm{FA}}$ of these techniques are evaluated over $350$ Monte-Carlo trials in Fig. 8 against SNR $\in[-25,25]~{}$dB.
@@
Figure 9: $P_{\mathrm{FA}}$ performance of KAN-10,5 histogram bins vs OSCFAR, with (a)-(d) represent input parameters of $P_{\mathrm{FA}}$ between $10^{-3}$ to $10^{-6}$.

### v3

@@ II Problem Statement
Consider a scenario with $K$ large targets with dimensions significantly larger than the signal wavelength. If the $k$th target is modeled with $I_{k}$ uncorrelated scatterers, the received signal is given by
@@
$\tau_{k,l}^{(i)}=\frac{2}{c}\big{(}R_{k}\pm\Delta{R_{k_{i}}}+(v_{k}\pm\delta v_{k_{i}})(l-1)T_{\mathrm{cri}}\big{)}$ (2)
@@ III-A Construction of the Binary hypothesis
It is well known that the down-converted receive signal $z(t,l)$ follows distinct distributions under each hypothesis: under $H_{0}$ (noise-only), it is modeled as $z(t,l)\sim\mathcal{CN}(0,\sigma^{2})$ ($\sigma$ is the standard deviation), while under $H_{1}$ (target present, modeled as Swerling-3), it follows $z(t,l)\sim\chi_{4}^{2}+\mathcal{CN}(0,\sigma^{2})$, $\chi_{4}^{2}$ denotes Chi-squared distribution of degree $4$. Post square-law detection, the RD cells $Z(f_{R},f_{D})$ under $H_{0}$ follow exponential distribution [24, 4]. However, under $H_{1}$, the distribution does not converge to a closed-form expression, necessitating a study of its empirical probability density function (pdf). Figure 2 presents histograms obtained over several min-max normalized RD segments. In Fig. 2(a), target presence causes most values to concentrate in the first bin, with minimal presence in higher bins. Fig. 2(b) shows a slower decay under the null hypothesis. While both histograms are observed to follow exponential distributions, Fig. 2(c) highlights their differing decay rates $\lambda_{0}\neq\lambda_{1}$, with the binary hypothesis as:
@@ IV-A Field data and Simulated data
The Medium Range Radar (MRR) specifications in Table I apply to both simulated and field data. Field data was collected using TI AWR2243 radar with a DCA1000 data capture card. A $4$m-long hatchback car is used for field acquisitions in scenarios that include the vehicle approaching and moving away with and without acceleration. For the specifications in Table I, the MRR provides a resolution of 0.35160.35160.3516~{}0.3516m and 0.30440.30440.3044~{}0.3044m///s in range and velocity, respectively. We construct RD segment with $17\times 7$ RD cells that correspond to (5.985.985.98~{}5.98m×$\times 2.13~{}$m///s). The Ground truth for real data is derived from manual annotation guided by the target’s GPS position.
Simulated data is generated by synthesizing a Swerling-3 target with $I_{k}\in[50,100]$ scatterers, distributed within $R_{k}\pm\Delta R_{k}$ (where $\Delta R_{k}\in[1.5,3]$ m), with scatterer powers drawn from a $\chi_{4}^{2}$ distribution. RCS values range from $19$ to $22~{}\mathrm{dBm}^{2}$ (side views), $8.7$ to $20.5~{}\mathrm{dBm}^{2}$(front views), and $14.4$ to $24.6~{}\mathrm{dBm}^{2}$ (rear views) [27], covering cars, SUVs, and trucks at random distances within MRR.
TABLE I: FMCW Radar Specification
$f_{0}$ FMCW slope $T_{cri}$ $f_{s}$ BW Samples Nchirps
@@ IV-B1 KAN for hypothesis classification
Unlike conventional evaluation of detection techniques, large-target detection has to be evaluated within the entire detection pipeline. Based on the number of histogram bins drawn from the RD segment, KAN is trained with MultKAN feature with specifications listed in Table II. KAN provides symbolic expressions at each output node in relation with the input. Detection accuracies presented in Table III are obtained by evaluating these symbolic expressions for (A) trained and tested on simulated data, (B) trained on simulated data, tested directly on field data, (C) Trained on simulated data, fine-tuned with 14 field samples, and tested on field data. The model handles deviations from (4) via transfer learning and can be extended to diverse kinds of targets.
TABLE II: KAN Model specifications and parameters
Width Grids spline Optimization,Loss Prune(Node:Edge)
[$M$,2] 3 3 LBFGS,Crossentropy 0.01:0.03
TABLE III: KAN symbolic expression performance
@@
From Table III, it is evident that the expressions derived from simulated data show moderate detection performance on field data (Scenario (B)). However, learning the KAN expression with as few as 14 samples in Scenario (C) significantly enhances detection performance, emphasizing the importance of transfer learning. The resulting symbolic expressions for $M=10$ and $M=5$ are obtained as
@@
Either side of the inequality serve as the test statistics for (5). One of these expressions is integrated with the RSP pipeline shown in Fig. 4. Notably, the KAN expression relies on only the first histogram bin $x_{0}$ for its decision in both expressions.
@@ IV-B2 Performance of detection pipeline
For the proposed method, $P_{D}$ is the percentage of target scattering response RD cells within the ground truth region. In OS-CFAR, a target is detected if at least one CUT within the ground truth region is identified. False alarms ($P_{\mathrm{FA}}$) are detected RD segments or CUTs with no corresponding target response. The proposed KAN-powered technique is compared with OS-CFAR with reference window matching the RD-segment and threshold design based on $P_{\mathrm{FA}}\in[10^{-3},10^{-6}]$ using both simulated and field data. For a fair comparison with OS-CFAR, the reference cells used for the threshold computation will be from the same RD segment excluding the CUT and the guard cells $[2,1]$ along range and Doppler axes, respectively. The $P_{\mathrm{D}}$ and $P_{\mathrm{FA}}$ of these techniques are evaluated over $350$ Monte-Carlo trials in Fig. 8 against SNR $\in[-25,25]~{}$dB.
@@
Figure 9: $P_{\mathrm{FA}}$ performance of KAN-10, 5 histogram bins vs OS-CFAR designed with $P_{\mathrm{FA}}$ between $10^{-3}$ to $10^{-6}$.